Status Report

NOAA GOES SATOPS Morning Report July 23, 2001

By SpaceRef Editor
July 23, 2001
Filed under , ,

GOES-M Launch Update

GOES-M was successfully launched this morning at 072301z. Apogee
was approximately 25
km low – overall, an excellent ride from the Atlas/Centaur.
Following separation, the sequence
of command activities leading up to solar array deployment were
successfully executed with a
maximum DoD of 22% on the batteries; well below the 60% maximum
allowable DoD and
consistent with GOES-9, GOES-10, and GOES-11 launches. The
partial solar array
deployment occurred without a hitch at ~08:39:15z. Initial
acquisitions at Diego Garcia, Perth,
and Canberra occurred close to predicts. Three-way doppler
between Diego Garcia and
Canberra was successfully processed by Flight Dynamics to aid in
quickly determining a
post-separation orbit.

Initial commanding was performed through Diego to initialize the
Attitude and Orbit Control
Electronics (AOCE #1), vent the propellant lines, perform 300
msec thruster firings using the
primary thrusters, command an off-axis sun acquisition, and
deploy the solar array outer panel.
All command sequences were performed flawlessly and the
spacecraft responded as expected.
Following solar array deployment, a command hand-over to Canberra
was performed to
supply Flight Dynamics sufficient data to determine a post-launch

Upcoming activities include the EPS/HEPAD Turn-on July 24 at
1443z and the 1st Apogee
Motor Firing (AMF #1) on July 25 at 0323z.


No change in the status of the POES spacecraft.

NOAA-14 rev 33798 / F at 1315z on July 21: The first two
scheduled recorder playbacks for
this pass started almost two minutes prior to the first command
time (FCT), and more than a
minute before first received telemetry (FRT). The result was the
loss of part of data sets
LAC1A (70%) and LAC2B (67.3%). Another recovery of these data was
not possible, due
to recorder scheduling limitations.

NOAA-15 rev 16570 / F at 1553z on July 21: – The first playback
from this pass started well
over a minute before FRT. (This is probably related to the use of
the FCDAS 13M ‘B’ system,
rather than the scheduled ‘C’ system, which was out of service.)
Only 86.8% of GAC5B was
received during this pass as a result. 100% of GAC5b was
recovered on rev 16571 / F at

NOAA-16 rev 4291 / F at 1633z on July 22: GAC1A received noisy
from spacecraft with
98.0% of the expected data recovered. No re-ingest required by

POES operations were nominal over the past 72 hours.


No change in the status of the DMSP spacecraft.

OPS48 rev 22164 R+09 FAIR/A at 18:09z on July 21: Cat III data
loss, 2885 s/c secs,
72.125 pbk secs. Playback was accidentally shut off during
Ephemeris Load. Recorder was
reset and all data played down to User, although User was unable
to process data.

DMSP operations were nominal over the past 72 hours.


No change in the status of the GOES spacecraft.

At 1152z on July 20, a GOES-10 frame break was observed and
attributed to high BER

At 2003z and 2041z on July 20, GOES-10 frame breaks were observed
and attributed to loss
of signal synchronization.

From 0441z to 0531z on July 21, the GOES-10 imager and sounder
were configured for
eclipse mode (non-imaging) as a result of a lunar shadow from
0449z to 0529z. Products
canceled included: the 0445z PACUS, 0452z Southern Hemi, 0500z
Northern Hemi, 0515z
PACUS, and 0530z Northern Hemi images and the 0501z CONUS and
0524z Northern
PACUS soundings. Total battery DoD was 6.79%.

At 0855z on July 22, Hurricane-rated antenna #1 (HR1) Transmitter
1 failed. WCDAS
personnel switched ASAP to Transmitter 2 and PDR and WEFAX was
restored at 0858z.
Commanding re-established at 0907z. The 0845z Full Disk image
lost multiple scans during
this time.

At 0822z on July 23, a GOES-8 frame break was observed and
attributed to loss of signal

GOES operations were nominal over the past 72 hours.

SpaceRef staff editor.